# Advanced Shaping of Quasi-Bessel Beams for High-Intensity Applications

### v1

@@ II Analytical description
To validate these theoretical predictions, we used the Axiprop optical propagation Python library [1], which relies on discrete Hankel and Fourier transforms. The simulations consider a beam of radius $R=50$ mm and an axiparabola mirror designed to produce a focal line of constant intensity and length $\delta=50$ mm, starting at $z=f_{0}=200$ mm (yielding, to first order, $r_{s}^{2}/R^{2}=(z-f_{0})/\delta$).
Figure 1 shows the intensity along the optical axis at the end (a) and at the beginning (b) of the focal line for a central hole radius $r_{h}=20$ mm. The agreement between the intensity predicted by Eq. (5) (orange curves) and the exact numerical result (blue curves) is excellent at the end of the focal line and fair at the beginning. This agreement degrades as the hole radius $r_{h}$ decreases, as illustrated in Fig. 1(c). This behavior occurs because the value of $\Psi^{\prime\prime}$ becomes smaller at low $r$, reducing the accuracy of the Taylor expansion in Eq. (2). Without a central hole, $\Psi$ varies too slowly for the stationary-phase approximation to hold, and the model breaks down. Nevertheless, the intensity profile still exhibits similar oscillations, as shown in Fig. 1(d), indicating that they do not originate from the hole but from edge effects associated with the sharp onset and termination of the focal line.
@@ III.1 Model
The numerical integration of Eq. (8) is plotted in Fig. 2, where it is compared to the results of an Axiprop simulation. The parameters used are identical to those in Sec. II, except for $r_{h}=10$ $m$, $L_{h}=4$ mm, and $n=100$. We observe an excellent agreement between the model and the simulations, as well as efficient suppression of the intensity of oscillations.
@@ III.2 Experiment
Figure 3 presents the recorded on-axis intensity profiles for both a sharp and a smooth hole, as defined by Eq.(7), for $r_{h}=11.6$ mm and $r_{h}=16.75$ mm, respectively, with $L_{h}=5.15$ mm. Each data point corresponds to the average of 10 images acquired with an exposure time of 200 ms. Similar to Figs. 1 and 2, pronounced oscillations are observed when using a sharp hole, while they are well suppressed with the Gaussian hole. A noticeable difference is observed in the intensity profile: instead of reaching the expected plateau after the hole, the intensity decreases. This behavior can be attributed to a trefoil aberration generated by the axiparabola, which distorts the focal intensity distribution. Introducing a trefoil aberration ($\Phi_{\text{trefoil}}(r,\theta)=2\pi\,a_{3}\,\cos\left(3\theta\left(\dfrac{r}{R}\right)^{3}\right)$ in Eq. (1) modifies the amplitude distribution by the factor
@@
Figure 3: Experimental on-axis intensity profiles obtained with a sharp hole (orange) and a Gaussian hole (blue) with $r_{h}=11.6$ mm and $r_{h}=16.75$ $m$ respectively. Theoretical prediction from simula is shown as dashed lines.
@@ IV.1 Model
This approach was validated through two Axiprop simulations performed for $r_{0}=10$ mm and $r_{0}=15$ mm, using the same axiparabola parameters as in Secs. II and III.1. As shown in Fig 4, oscillations are strongly damped for $r_{0}=10$ mm, similar to amplitude shaping, and almost completely suppressed for $r_{0}=15$ mm. Since Eq. (6) is satisfied, the on-axis intensity can be computed directly from Eq. (4). The resulting profiles, plotted in Fig. 4, show excellent agreement with the numerical simulations. Furthermore, the intensity evolution within the rising ramp can be obtained analytically by deriving $r_{s}(z)$ for the phase defined in Eq. (10), leading to
@@ IV.2 Experiment
Figure 5 shows the recorded intensity profiles for the unmodulated beam and for the beam shaped with a phase modulation of $r_{0}=10$ mm. Each data point corresponds to the average of 20 images acquired with an exposure time of 100 ms. We observe that the oscillations are strongly damped, with a long rising ramp, in good agreement with the simulations.
@@ V.1 Resolution limit
Interestingly, the achievable resolution is independent of the position along the focal line and depends only on the numerical aperture and the focal depth. To validate this estimate, we performed a set of Axiprop simulations. We considered a radial intensity profile featuring two Gaussian holes located at $r_{1}=0.45R$ and $r_{2}=0.9R$, described by $E_{i}(r)=E_{i0}(r)-E_{0}\exp\left(-(r-r_{i})^{2}/L_{ri}^{2}\right)$ where $i\in\{1,2\}$, $E_{i0}(r)$ is the radial profile without a hole, and $L_{ri}=L_{r}(r_{i})$. Figure 6 presents the simulation results for $R=50$ mm, $f_{0}\in\{200,400\}$ $m$, $\delta\in\{30,60\}$ mm and various values of $L_{r}\in\{1,3,6,9\}L_{r_{0}}$.
Figure 6: On-axis intensity profiles obtained from simulations for different axiparabola parameters. (a) $f_{0}=200$ mm and $\delta=30$ $m$. (b) $f_{0}=200$ mm and $\delta=60$ $m$. (c) $f_{0}=400$ mm and $\delta=60$ mm. The curves correspond to different values of $L_{r}\in\{1,3,6,9\}L_{r_{0}}$

### v2

@@ II Analytical description
To validate these theoretical predictions, we used the Axiprop optical propagation Python library [1], which relies on discrete Hankel and Fourier transforms. The simulations consider a beam of radius $R=50$ mm and an axiparabola mirror designed to produce a focal line of constant intensity and length $\delta=50$ $m$, starting at $z=f_{0}=200$ $m$ (yielding, to first order, $r_{s}^{2}/R^{2}=(z-f_{0})/\delta$).
Figure 1 shows the intensity along the optical axis at the end (a) and at the beginning (b) of the focal line for a central hole radius $r_{h}=20$ $m$. The agreement between the intensity predicted by Eq. (5) (orange curves) and the exact numerical result (blue curves) is excellent at the end of the focal line and fair at the beginning. This agreement degrades as the hole radius $r_{h}$ decreases, as illustrated in Fig. 1(c). This behavior occurs because the value of $\Psi^{\prime\prime}$ becomes smaller at low $r$, reducing the accuracy of the Taylor expansion in Eq. (2). Without a central hole, $\Psi$ varies too slowly for the stationary-phase approximation to hold, and the model breaks down. Nevertheless, the intensity profile still exhibits similar oscillations, as shown in Fig. 1(d), indicating that they do not originate from the hole but from edge effects associated with the sharp onset and termination of the focal line.
@@ III.1 Model
The numerical integration of Eq. (8) is plotted in Fig. 2, where it is compared to the results of an Axiprop simulation. The parameters used are identical to those in Sec. II, except for $r_{h}=10$ $m$, $L_{h}=4$ $m$, and $n=100$. We observe an excellent agreement between the model and the simulations, as well as efficient suppression of the intensity of oscillations.
@@ III.2 Experiment
Figure 3 presents the recorded on-axis intensity profiles for both a sharp and a smooth hole, as defined by Eq.(7), for $r_{h}=11.6$ $m$ and $r_{h}=16.75$ $m$, respectively, with $L_{h}=5.15$ $m$. Each data point corresponds to the average of 10 images acquired with an exposure time of 200 ms. Similar to Figs. 1 and 2, pronounced oscillations are observed when using a sharp hole, while they are well suppressed with the Gaussian hole. A noticeable difference is observed in the intensity profile: instead of reaching the expected plateau after the hole, the intensity decreases. This behavior can be attributed to a trefoil aberration generated by the axiparabola, which distorts the focal intensity distribution. Introducing a trefoil aberration ($\Phi_{\text{trefoil}}(r,\theta)=2\pi\,a_{3}\,\cos\left(3\theta\left(\dfrac{r}{R}\right)^{3}\right)$ in Eq. (1) modifies the amplitude distribution by the factor
@@
Figure 3: Experimental on-axis intensity profiles obtained with a sharp hole (orange) and a Gaussian hole (blue) with $r_{h}=11.6$ $m$ and $r_{h}=16.75$ $m$ respectively. Theoretical prediction from simulation is shown as dashed lines.
@@ IV.1 Model
This approach was validated through two Axiprop simulations performed for $r_{0}=10$ $m$ and $r_{0}=15$ $m$, using the same axiparabola parameters as in Secs. II and III.1. As shown in Fig 4, oscillations are strongly damped for $r_{0}=10$ $m$, similar to amplitude shaping, and almost completely suppressed for $r_{0}=15$ $m$. Since Eq. (6) is satisfied, the on-axis intensity can be computed directly from Eq. (4). The resulting profiles, plotted in Fig. 4, show excellent agreement with the numerical simulations. Furthermore, the intensity evolution within the rising ramp can be obtained analytically by deriving $r_{s}(z)$ for the phase defined in Eq. (10), leading to
@@ IV.2 Experiment
Figure 5 shows the recorded intensity profiles for the unmodulated beam and for the beam shaped with a phase modulation of $r_{0}=10$ $m$. Each data point corresponds to the average of 20 images acquired with an exposure time of 100 ms. We observe that the oscillations are strongly damped, with a long rising ramp, in good agreement with the simulations.
@@ V.1 Resolution limit
Interestingly, the achievable resolution is independent of the position along the focal line and depends only on the numerical aperture and the focal depth. To validate this estimate, we performed a set of Axiprop simulations. We considered a radial intensity profile featuring two Gaussian holes located at $r_{1}=0.45R$ and $r_{2}=0.9R$, described by $E_{i}(r)=E_{i0}(r)-E_{0}\exp\left(-(r-r_{i})^{2}/L_{ri}^{2}\right)$ where $i\in\{1,2\}$, $E_{i0}(r)$ is the radial profile without a hole, and $L_{ri}=L_{r}(r_{i})$. Figure 6 presents the simulation results for $R=50$ mm, $f_{0}\in\{200,400\}$ $m$, $\delta\in\{30,60\}$ $m$ and various values of $L_{r}\in\{1,3,6,9\}L_{r_{0}}$.
Figure 6: On-axis intensity profiles obtained from simulations for different axiparabola parameters. (a) $f_{0}=200$ $m$ and $\delta=30$ $m$. (b) $f_{0}=200$ $m$ and $\delta=60$ $m$. (c) $f_{0}=400$ $m$ and $\delta=60$ $m$. The curves correspond to different values of $L_{r}\in\{1,3,6,9\}L_{r_{0}}$